OpenStudy (anonymous):

Solve log11 (y + 8) + log11 4 = log11 60

OpenStudy (breanneh):

log11 (y + 8) + log11 4 = log11 60 it can rewriten be log11 (y + 8)4 = log11 60 now, cancel out the log 11 to both sides, giving u (y+8)4 = 60 solve for y

OpenStudy (breanneh):

So it would equal 7.

OpenStudy (anonymous):

thanks you bunches!!!

OpenStudy (breanneh):

Sure thing!